# Dual-Unitary Classical Shadow Tomography

### v1

@@ II.7 Ramifications for the shadow norm and prediction
To summarize, we’ve introduced dual-unitary circuit shadow tomography as a more feasible alternative to shadow tomography based on Clifford circuits. The feasibility derives from the simple implementation of two-local dual unitary gates from standard logical operations and local scrambling. Compared to brick-wall Clifford circuits, dual unitaries generally perform worse because they more quickly disperse the Pauli weight of a fixed operator with size $k<n$. Despite this inherent weakness for predicting small operators, dual-unitary gates more quickly relax the Pauli weight and can approach the Clifford measurement limit within the bulk faster. Hence, dual-unitary circuits are advantageous for predicting operators with full or almost support $k\approx n$, because in that case the Pauli weight is already dispersed. The advantage can be observed in Fig. 14. We characterize the advantage in a $n$-independent way by studying the scaling exponent $\beta$ of the shadow norm of a $n$-local Pauli string. We find that this scaling exponent, and therefore the overall sample complexity, is lower for dual-unitary gates with large enough $\alpha$. Furthermore, we find that the dual-unitary gates more quickly approach the Clifford limit $\beta=2$. Combined with previous results from Akhtar et al. (2023), this suggests a more efficient method for the prediction of general operators with full support.
@@ III Conclusions
In this work, we explore the average Pauli weight dynamics in dual-unitary brick-wall circuits. We find that it can be described in terms of chiral quasi-particles that propagate at fixed velocity. These particles can fuse into each other and emit one another with a probability determined by the dual-unitary parameter $\alpha$. In particular, at the swap point, $\alpha$ vanishes and the different movers do not interact. We analyze the behavior of the occupation $\rho(x,t)$ as a function of $x,t$ and $\alpha$ and find that for initially contiguous operators it exhibits a ballistic front that moves with maximal light-cone velocity $v_{LC}=1$, and rapidly thermalizes the interior of the light-cone. One advantage of this is that finite depth dual-unitary circuits rapidly approximate the Clifford measurement limit of classical shadow tomography within the bulk, suggesting an advantage for predicting extensive operators.
@@
Our work focused on qubit systems for simplicity and relevance to classical shadow tomography, but in principle one could study how varying local Hilbert space dimension affects the results and observations of this paper. Furthermore, the particular brick-wall geometry was also chosen with practical applications in mind, but in principle one could consider different arrangements of brick-wall gates. Since the brick-wall structure seems essential for the emergence of the chiral quasiparticles, one can ask which of the features of the dual-unitary Pauli weight dynamics are independent of geometry.
@@
In principle, the results of this paper suggest that finite depth, dual-unitary classical shadow tomography could be a powerful tool for predicting extensive operators. We would like for this intuition to be demonstrated on an actual quantum device. Furthermore, it would be helpful to evaluate whether such circuits would be better at fidelity estimation than finite depth Clifford circuits, which were shown to be a more practical alternative to Clifford measurements for predicting low rank operators Akhtar et al. (2023).
@@
In this work, we use a mean-field treatment instead of an exact treatment of the Pauli weight dynamics. We find that for $\rho(x,t)$, there is practically no difference. However, for calculating the full shadow norm, it is possible that neglecting entanglement for the Pauli weight state could create some discrepancy in our estimation for e.g. the shadow norm of observables with a more complicated spatial support. Therefore, a fuller treatment of the Pauli weight dynamics could be useful.
@@
The emergence of chiral excitations reminiscent of relativistic QFTs in dual-unitary Pauli weight is interesting from the perspective of AdS/CFT correspondence, according to which the entanglement of conformal field theories is related to the geometry of Anti de-Sitter space. Since the Pauli weight is closely related to the second Renyi entropy of the quantum circuit, and since dual-unitary circuits can mimic the entanglement structure in conformal field theories, it stands to reason the emergence of relativity in the Pauli weight dynamics could be a more profound consequence of this correspondence. Can we relate the Pauli weight dynamics to geometry of AdS space-time? In what other ways does relativity manifest in these circuits?
@@ IV Acknowledgements
A.A.A. and Y.Z.Y. are supported by a startup fund from UCSD. N.A. is a KBR employee working under the Prime Contract No. 80ARC020D0010 with the NASA Ames Research Center. J.M. is thankful for support from NASA Academic Mission Services, Contract No. NNA16BD14C. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under the contract No. DE-AC02-07CH11359; with this work funded under the NASA-DOE interagency agreement SAA2-403602 governing NASA’s work as part of the SQMS center. The United States Government retains, and by accepting the article for publication, the publisher acknowledges that the United States Government retains, a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work, or allow others to do so, for United States Government purposes.

### v2

@@ II.7 Ramifications for the shadow norm and prediction
To summarize, we’ve introduced dual-unitary circuit shadow tomography as a more feasible alternative to shadow tomography based on Clifford circuits. The feasibility derives from the simple implementation of two-local dual unitary gates from standard logical operations and local scrambling. Compared to brick-wall Clifford circuits, dual unitaries generally perform worse because they more quickly disperse the Pauli weight of a fixed operator with size $k<n$. Despite this inherent weakness for predicting small operators, dual-unitary gates more quickly relax the Pauli weight and can approach the Clifford measurement limit within the bulk faster. Hence, dual-unitary circuits are advantageous for predicting operators with full or almost support $k\approx n$, because in that case the Pauli weight is already dispersed. The advantage can be observed in Fig. 14. We characterize the advantage in a $n$-independent way by studying the scaling exponent $\beta$ of the shadow norm of a $n$-local Pauli string. We find that this scaling exponent, and therefore the overall sample complexity, is lower for dual-unitary gates with large enough $\alpha$. Furthermore, we find that the dual-unitary gates more quickly approach the Clifford limit $\beta=2$. Combined with previous results from [48], this suggests a more efficient method for the prediction of general operators with full support. Furthermore, we find that for small $n$, the advantage is actually amplified due to the finite size effects, which can be seen in Fig. 14 c.
Figure 15: We study the effects of the boundary on the average occupation $\rho(x,t)$ in the Clifford and dual-unitary cases, using numerics from a $n=24$ qubit system for circuit depths $t=2,4,6$, as an illustrative example. (a). In the Clifford case, we find a persistent density of occupation on the boundaries of the system. Because there is no notion of left or right movers in this case, the effect of the boundary is to simply prevent weight from “leaking” out, resulting in a higher probability of having operators with support on the boundary. This persistent density of occupation on the boundaries increases the shadow norm according to Eq. (63). In short, the boundary in the brick-wall Clifford case traps excitations, thereby increasing the shadow norm. (b). In the dual-unitary case, there is no persistent density of boundary excitations. This is because excitations in the dual-unitary case are chiral i.e. either left-moving or right-moving. These excitations reflect off the boundary and propagate into the bulk, where the fast-thermalizing nature of the dual-unitary gates causes them to quickly dissipate. Because the boundaries don’t trap excitations in the dual-unitary case, the finite size effects are expected to be favorable to the primary claim of this paper. It is also worth noting that the fast-thermalization properties of dual-unitary circuits manifests in the scale of the $y$-axis of the figures decaying much faster for the dual-unitary case when compared with the Clifford case.
In the preceding analysis, we’ve largely ignored the boundary and finite-size effects and correspondingly used the mean-field approximation, which assumes a large system. An important concern could be that such circuits may not be advantageous in near-term devices, where large numbers of qubits cannot be simulated. We now explain why: the boundary effects are in favor of the primary claim of this paper, that the dual-unitary brick-wall circuits are advantageous compared to their Clifford counterparts for predicting operators with extensive support. To understand why, let’s explore the role of the boundary in both cases. In the Clifford case, the boundary amounts to the absence of a transfer matrix that could “leak” weight outside the system. Because there is no notion of chirality in the Clifford case, when an excitation in the occupation $\rho(x,t)$ reaches the boundary, it simply sits there without diffusing. The build-up of occupation density on the boundaries is actually disadvantageous for prediction, since it slows down the overall thermalization of the system which brings the protocol closer to the global Clifford limit which is known to be advantageous for extensive operators.
Contrarily, in the dual-unitary case, the excitations must be either left or right-moving. This means that when an excitation in the occupation meets the boundary, it reflects and propagates back into the bulk. Once it enters the bulk, the bulk thermalizing properties of the dual-unitary, which are studied in the mean-field approximation, become dominant. In particular, because the dual-unitaries saturate more quickly in the bulk, the reflected excitations quickly dissipate and the overall weight of the system approaches the global Clifford limit more quickly. Unlike the Clifford case, the boundary excitations, which increase the shadow norm, are forced back into the bulk. This effect is unique to dual-unitary circuits, where there is a notion of chirality, and explains why the boundary does not undermine the primary result of this paper.
We numerically study these boundary effects for a $n=24$-qubit system for shallow circuit depths of $t=2,4,6$ in Figure 15. These numerical simulations highlight a perhaps nontrivial fact that dual-unitary shadow tomography schemes are just as useful for system sizes accessible in the near-term (for operators with full support). Said another way, our scheme is robust to finite-size effects.
@@ III Conclusions
In this work, we explore the average Pauli weight dynamics in dual-unitary brick-wall circuits. We find that it can be described in terms of chiral quasi-particles that propagate at fixed velocities. These particles can fuse into each other and emit one another with a probability determined by the dual-unitary parameter $\alpha$. In particular, at the swap point, $\alpha$ vanishes and the different movers do not interact. We analyze the behavior of the occupation $\rho(x,t)$ as a function of $x,t$ and $\alpha$ and find that for initially contiguous operators it exhibits a ballistic front that moves with maximal light-cone velocity $v_{LC}=1$, and rapidly thermalizes the interior of the light-cone. One advantage of this is that finite depth dual-unitary circuits rapidly approximate the Clifford measurement limit of classical shadow tomography within the bulk, suggesting an advantage for predicting extensive operators.
@@
We focused on qubit systems for simplicity and relevance to classical shadow tomography, but in principle, one could study how varying local Hilbert space dimension affects the results and observations of this paper. Furthermore, the particular brick-wall geometry was also chosen with practical applications in mind, but in principle, one could consider different arrangements of brick-wall gates. Since the brick-wall structure seems essential for the emergence of the chiral quasiparticles, one can ask which of the features of the dual-unitary Pauli weight dynamics are independent of geometry.
@@
In principle, the results of this paper suggest that finite depth, dual-unitary classical shadow tomography could be a powerful tool for predicting extensive operators. We would like for this intuition to be demonstrated on an actual quantum device. Furthermore, it would be helpful to evaluate whether such circuits would be better at fidelity estimation than finite depth Clifford circuits, which were shown to be a more practical alternative to Clifford measurements for predicting low-rank operators [48].
@@
We use a mean-field treatment instead of an exact treatment of the Pauli weight dynamics. We find that for $\rho(x,t)$, there is practically no difference. However, for calculating the full shadow norm, it is possible that neglecting entanglement for the Pauli weight state could create some discrepancy in our estimation for e.g. the shadow norm of observables with a more complicated spatial support. Therefore, a fuller treatment of the Pauli weight dynamics could be useful.
@@
The emergence of chiral excitations reminiscent of relativistic QFTs in dual-unitary Pauli weight is interesting from the perspective of AdS/CFT correspondence, according to which the entanglement of conformal field theories is related to the geometry of Anti de-Sitter space. Since the Pauli weight is closely related to the second Renyi entropy of the quantum circuit, and since dual-unitary circuits can mimic the entanglement structure in conformal field theories, it stands to reason the emergence of relativity in the Pauli weight dynamics could be a more profound consequence of this correspondence. Can we relate the Pauli weight dynamics to the geometry of AdS space-time? In what other ways does relativity manifest in these circuits?
@@ IV Acknowledgements
A.A.A. and Y.Z.Y. are supported by a startup fund from UCSD and the NSF Grant No. DMR-2238360. N.A. is a KBR employee working under the Prime Contract No. 80ARC020D0010 with the NASA Ames Research Center. J.M. is thankful for support from NASA Academic Mission Services, Contract No. NNA16BD14C. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359; with this work funded under the NASA-DOE interagency agreement SAA2-403602 governing NASA’s work as part of the SQMS center. The United States Government retains, and by accepting the article for publication, the publisher acknowledges that the United States Government retains, a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this work or allow others to do so, for United States Government purposes.